# SocialJax: An Evaluation Suite for Multi-agent Reinforcement Learning in Sequential Social Dilemmas

### v1

@@ 3.3 Algorithms in JAX
Due to IPPO’s widespread adoption and stability (Schulman et al., 2017; De Witt et al., 2020), we use it as our baseline algorithm. IPPO is also adopted by PureJaxRL (Lu et al., 2022) and JaxMARL (Rutherford et al., 2023). We leverage convolutional neural networks to process state information from the SocialJax environments. To encourage prosocial behavior, we introduce common rewards, while individual rewards are used to incentivize selfishness in agents.
@@
Common Rewards refers to a scenario where all agents in a multi-agent system share a single, unified reward signal derived from the environment. This approach ensures that all agents are aligned towards achieving the same objective, promoting collaboration and coordination among them. By sharing a common reward, agents are incentivized to work together and make decisions that benefit the collective, rather than focusing solely on individual gains. This can help prevent conflicts and encourage cooperative behavior.
@@
Individual Rewards indicate that each agent is assigned its own reward, inherently encouraging selfish behavior. This approach allows agents to prioritize maximizing their personal success over collective goals. We examine the performance of selfish agents in social dilemma environments and compare it against their performance under a common reward scheme.
@@
Territory $2.5\times 10^{3}$ $6.6\times 10^{2}$ $8.1\times 10^{4}$ $4.6\times 10^{5}$ $9.5\times 10^{5}$
@@
4 Experiments
4.1 Setup
To validate our library, we perform thorough tests of both Common Reward IPPO and Individual Reward IPPO algorithms across a diverse range of SocialJax environments. In addition to these experiments, we plot Schelling diagrams to further verify that our environments effectively represent social dilemmas. This evaluation ensures that the environments within SocialJax are suitable for studying sequential social dilemmas.
We assess the performance of our JAX-based environment on GPU and compare its speed with the implementations provided by MeltingPot 2.0 (Agapiou et al., 2022). When taking random actions, the speed of SocialJax is approximately 70-400 times faster than Stable-Baselines3 and 40-60 times faster than RLlib, both open-sourced by MeltingPot 2.0.
@@ 4.2 Results
Speed Benchmarks From Table 1, we observe a significant increase in the speed of the training pipeline with the addition of more parallel environments. Since both the environments and algorithms are vectorized using JAX, they are well-suited for GPU acceleration. In Table 1, we tested up to 4096 environments in parallel.
We compare the performance of SocialJax and MeltingPot 2.0 (Agapiou et al., 2022) by measuring steps per second on the same hardware, which consists of an NVIDIA A100 GPU and 14 CPU cores. As shown in Table 1, we evaluate the speed of SocialJax across various environments and environment counts using random actions. We observe a significant speedup when running environments in parallel on a GPU compared to running a single environment. The original environments run on a single CPU, and achieving the parallel efficiency of SocialJax would require scaling to hundreds of CPUs. We also compare the performance of our training pipeline with the training speed of the Stable-Baselines3 and RLlib implementations with TensorFlow, based on the publicly available code from Melting Pot 2.0. Note that the experimental results in Melting Pot 2.0 (Agapiou et al., 2022) cannot be directly compared, as the code related to their reported results has not been released. The open-source versions of the algorithms in Melting Pot 2.0 are based on RLlib and Stable Baselines3. However, the versions of Actor-Critic Baseline (ACB) (Espeholt et al., 2018), V-MPO (Song et al., 2019), and OPRE (Vezhnevets et al., 2020) used in their reports have not been made available as open source.
Additionally, since MeltingPot 2.0 is pixel-based environments while our SocialJax environments are grid-based, it is not feasible to align the environment and neural network settings directly. Therefore, we only perform real-time comparisons for training scenarios where convergence is achieved. For relatively simple environments such as Coins, SocialJax can complete 1e9 timesteps within 3 hours, whereas training with Melting Pot 2.0 using Stable-Baselines3 requires approximately 1,300 hours (on an NVIDIA A100 GPU with 14 CPU cores), training with Melting Pot 2.0 using RLlib takes around 150 hours. Consequently, conducting social dilemma experiments with Melting Pot 2.0 is highly challenging without a large amount of CPU resources. Training Coins in SocialJax is about 50-400 times faster compared to Coins in Melting Pot 2.0. For relatively complex environments such as Clean Up and Territory, SocialJax is approximately 50-140 times faster than Melting Pot 2.0.
@@
Algorithm Performance We compare the performance of common reward IPPO and individual reward IPPO across different environments. In the default individual reward setting, each agent receives a reward based on the joint-action, independent of the payoffs of other agents. The environments use homogeneous reward functions for all agents. This creates incentives for agents to free-ride, or to over exploit common pool resources, because only they themselves benefit from their actions, while the social costs are shared by all.
@@
In contrast, a common reward, or team reward means that all agents share a single reward signal, typically the total payoff of all agents in the environment. The common reward encourages cooperation, as the objective of each agent is to maximize the total reward, making individual agents more inclined to adopt collaborative behaviors. In environments such as Coins (Figure 2(a)), Harvest (Figure 2(b), Figure 2(c)), Clean Up (Figure 2(d)), and Coop mining (Figure 2(e)), a common reward typically performs better than individual rewards, as the agents attempt to maximize the total payoff, removing the incentives to engage in selfish behaviors.
However, using a common reward structure can exacerbate the credit assignment problem. Specifically, if an agent takes an arbitrary action concurrently with a teammate who performs a successful action generating a reward, the agent may mistakenly attribute that reward to its own action. This misattribution hinders accurate learning and policy improvement, leading to suboptimal coordination. We observe this phenomenon explicitly in our Territory environment (see Figure 2(f)), where common reward agents fail to identify effective collaborative policies. Consequently, their overall performance remains inferior to agents trained with individual rewards, reflecting a clearer assignment of credit.
@@ 4.3 Social Dilemma Environments Attributes
Coins (Figure 3(a)), Harvest: Open (Figure 3(b)), Harvest: Closed (Figure 3(c)) exhibit fear: namely that agents would prefer to defect if some of their co-players are defecting. An agent actually does better cooperating if all their co-players are cooperating. In Harvest, we observe that it only takes a single defector agent to have a severely negative impact on the returns of other players. This is because defector agents will typically harvest the last apple in a patch, permanently depleting it. The cooperator policies have learned to leave a greater amount of apples in each patch, which causes a higher regrowth rate, so they generally exhibit a greater tendency to abstain from harvesting in the presence of a defector.
In Clean Up (Figure 3(d)), we instead observe greed, as there is an incentive to defect if a sufficient number of co-players are cooperating. This is due to that fact that cooperators will clean the pollution, whereas the defector will not, but it requires a minimum of three cooperators to clean the river for apples to grow. Coop Mining (Figure 3(e)) primarily exhibits the greed property. In this environment, if a sufficient number of agents choose to cooperate by focusing on collecting gold, other agents can obtain a higher payoff by defecting. This is because the cooperators tend to neglect the ore, allowing the defector to collect more ore.
Due to the difficulties of credit assignment in Territory (Figure 3(f)), IPPO with a common rewards fails to learn effective policies: they perform consistently worse than the policies with individual rewards. We believe that this environment should be a social dilemma, because there are personal incentives that undermine the collective welfare, but more advanced algorithms or training methodologies are required to find effective cooperative policies.

### v2

@@ 3.3 Algorithms in JAX
Due to IPPO’s widespread adoption and stability Schulman et al. (2017); De Witt et al. (2020), we implement IPPO under both parameter-sharing and non-parameter-sharing settings. The parameter-sharing variant of IPPO leads to faster training by using a shared policy neural network across all agents. The non-parameter-sharing variant, although slower, preserves agent-specific policies that prevent convergence to uniform conventions. The IPPO training curves shown in Figure 2 are based on the parameter sharing variant. Additional comparisons between parameter sharing and non-parameter sharing IPPO can be found in the Appendix B.1. To encourage prosocial behavior, we introduce common rewards, while individual rewards are used to incentivize selfishness in agents. Additionally, we include MAPPO Yu et al. (2022) as a representative centralized learning algorithm for MARL. Finally, we implement SVO McKee et al. (2020) as a representative method specifically designed to handle social dilemma scenarios, where agents can balance individual and collective interests.
@@
IPPO Common Rewards refers to a scenario where all agents in a multi-agent system share a single, unified reward signal derived from the environment. This approach ensures that all agents are aligned towards achieving the same objective, promoting collaboration and coordination among them. By sharing a common reward, agents are incentivized to work together and make decisions that benefit the collective, rather than focusing solely on individual gains. This can help prevent conflicts and encourage cooperative behavior.
@@
IPPO Individual Rewards indicates that each agent is assigned its own reward, inherently encouraging selfish behavior. This approach allows agents to prioritize maximizing their personal success over collective goals.
MAPPO We implement Multi-Agent PPO (MAPPO) Yu et al. (2022) using both the PureJaxRL Lu et al. (2022) and JaxMARL Rutherford et al. (2023) frameworks. As a centralized learning approach, MAPPO uses the centralized value function to help incentivize agents to perform prosocial behaviors.
SVO We implement SVO McKee et al. (2020) as an intrinsic motivation framework to encourage cooperative behavior in our environments. To capture the relative reward distribution, reward angle $\theta(R)=\arctan\left(\frac{\bar{r}_{-i}}{r_{i}}\right)$ is defined by SVO, where $r_{i}$ is the reward received by agent $i$, and $\bar{r}_{-i}=\frac{1}{n-1}\sum_{j\neq i}r_{j}$ is the average reward of all other agents. Each agent is assigned a target SVO angle $\theta^{\text{SVO}}$ that reflects its desired distribution of reward. The agent’s utility function combines extrinsic and intrinsic rewards as
$U_{i}(s,o_{i},a_{i})=r_{i}-w\cdot|\theta^{\text{SVO}}-\theta(R)|$
where $w$ controls the strength of the Social Value Orientation. We restrict SVO preferences to the non-negative quadrant, i.e., $\theta\in[0^{\circ},90^{\circ}]$, where $\theta=90^{\circ}$ represents an Altruistic setting and $\theta=0^{\circ}$ represents an Individualistic setting.
3.4 Metrics for SocialJax Environments
Since an agent’s behavior in the environments cannot be fully captured by the return alone, we devised a tailored metric for each environment to quantify whether agents tend toward cooperation or competition. We define the semantics of the metrics for each environment as follows:
Coins: Since each agent collecting their own-color coin does not harm the other, whereas collecting the other agent’s coin imposes a penalty on them, we evaluate cooperation by counting how many coins each agent collects that match their assigned color.
Commons Harvest: In this environment, agents must maintain a sustainable number of apples to ensure they continue to respawn. Therefore, we measure the number of apples remaining on the map to assess whether agents are capable of preserving the long-term collective interest.
Clean Up: Agents need to clean polluted river tiles to enable apple regrowth. We evaluate their cooperative or selfish behavior by counting the number of cleaned water tiles.
Coop Mining: In this environment, agents can obtain higher rewards by cooperating to mine gold, while those who choose not to cooperate can only mine iron for a smaller reward. Therefore, we evaluate agents’ cooperative behavior by measuring the amount of gold mined.
Mushrooms: The environment contains multiple types of mushrooms. Consuming blue mushrooms does not benefit the agent itself but provides rewards to the other agents. We thus measure the number of blue mushrooms consumed to evaluate the agent’s willingness to sacrifice for the common good.
Gift Refinement: In this environment, agents can choose to gift collected tokens to others, which results in receiving a larger reward in return. We thus count the number of received tokens to evaluate whether agents are willing to trust each other.
Prisoners Dilemma: Arena: The environment contains two types of resources, cooperative and competitive, and we introduce the quantity of cooperative resources collected by agents as a new metric.
4 Experiments
4.1 Setup
To validate our library, we perform thorough tests of Common Reward IPPO, Individual Reward IPPO, MAPPO, and SVO algorithms across a diverse range of Sequential Social Dilemma environments and analyze the returns and designed cooperation metrics. We assess the performance of our JAX-based environment on a GPU and compare its speed with the implementations provided by MeltingPot 2.0 Agapiou et al. (2022). In addition to these experiments, we plot Schelling diagrams to further verify that our environments effectively represent social dilemmas.
@@ 4.1 Setup
Harvest: Partnership $3.7\times 10^{3}$ $1.2\times 10^{3}$ $1.1\times 10^{5}$ $5.0\times 10^{5}$ $7.9\times 10^{5}$
@@
Mushrooms $4.2\times 10^{3}$ $1.4\times 10^{3}$ $1.4\times 10^{5}$ $6.3\times 10^{5}$ $9.8\times 10^{5}$
Gift Refinement $4.1\times 10^{3}$ $1.5\times 10^{3}$ $1.5\times 10^{5}$ $7.0\times 10^{5}$ $1.2\times 10^{6}$
Prisoners Dilemma: Arena $4.5\times 10^{3}$ $2.2\times 10^{3}$ $2.3\times 10^{5}$ $1.3\times 10^{6}$ $2.7\times 10^{6}$
@@ 4.2 Results
Speed Benchmarks Since both the environments and algorithms are vectorized using JAX, they are well suited for GPU acceleration. In Table 1, we test up to 4096 environments in parallel and observe a significant increase in the speed of the training pipeline with the addition of more parallel environments.
We compare the performance of SocialJax and MeltingPot 2.0 Agapiou et al. (2022) by measuring steps per second on the same hardware, which consists of an NVIDIA A100 GPU and 14 CPU cores. As shown in Table 1, we evaluate the speed of SocialJax across various environments and environment counts using random actions. We observe a significant speedup when running environments in parallel on a GPU compared to running a single environment. To match the parallel performance of SocialJax, the original environments would need to scale up across hundreds of CPUs. Regarding wall clock time, SocialJax is capable of completing 1e9 timesteps within 3 hours for relatively simple environments such as Coins, whereas training with Melting Pot 2.0 using Stable-Baselines3 and RLlib takes around 1,300 hours and 150 hours, respectively. Consequently, conducting social dilemma experiments with Melting Pot 2.0 is highly challenging without a large amount of CPU resources. Training Coins in SocialJax is about 50-400 times faster compared to Coins in Melting Pot 2.0. For relatively complex environments such as Clean Up, SocialJax is approximately 50-140 times faster than the original environments. There are more details about the wall clock time comparison in the Appendix B.2.
(a) Coins
(b) Commons Harvest: Open
(c) Commons Harvest: Closed
(d) Commons Harvest: Partnership
(e) Clean Up
(f) Coop Mining
(g) Mushrooms
(h) Gift Refinement
(i) Prisoners Dilemma: Arena
Figure 2: Training curves for a range of SocialJax environments. IPPO (shared parameters) with Common Reward encourages collective interests, leading to higher overall returns, Individual Reward primarily drives selfish behavior, often resulting in lower returns. MAPPO is included as a centralized baseline. The SVO curve represents the case with a social angle $\theta=90^{\circ}$.
@@
IPPO Common reward and IPPO Individual reward We compare the performance of common reward IPPO (IPPO-CR) and individual reward IPPO (IPPO-IR) across different environments. In the default individual reward setting, each agent receives a reward based on the joint-action, independent of the payoffs of other agents. This creates incentives for agents to over exploit common pool resources, because only they themselves benefit from their actions, while the social costs are shared by all.
@@
In contrast, a common reward, or team reward means that all agents share a single reward signal, typically the total payoff of all agents in the environment. The common reward encourages cooperation, as the objective of each agent is to maximize the total reward, making individual agents more inclined to adopt collaborative behaviors. In Figure 2, a common reward typically performs better than individual rewards, as the agents attempt to maximize the total payoff.
MAPPO As a centralized learning method, MAPPO has access to information from all agents, which leads to strong performance in Clean Up (Figure 2(e)), and Mushrooms (Figure 2(g)). However, MAPPO exhibits less stable training dynamics, as its centralized critic aggregates observations from all agents, increasing the learning difficulty of the network. In particular, it struggles with Commons Harvest: Partnership (Figure 2(d)) and Gift Refinement (Figure 2(h)), where it significantly under performs IPPO-CR.
SVO (McKee et al., 2020) provides a powerful framework for exploring agents’ cooperative and competitive tendencies. We apply SVO across all SocialJax environments to investigate how the preference angle $\theta$ influences agents’ returns. We first fix $\theta=90^{\circ}$ and sweep over different weight values $w$ to identify the optimal $w$. Then, holding the weight fixed, we sweep over $\theta$ to examine how agent behavior and returns vary with the angle. As summarized in Table 2, the collective return increases with larger values of $\theta$, as agents become more inclined to cooperate at higher angles. When $\theta=0^{\circ}$ (Individualistic orientation), agents achieve the lowest returns, whereas at $\theta=90^{\circ}$ (Altruistic orientation), they attain the highest returns.
IPPO with a common reward exhibits greater stability than both MAPPO and SVO, making it a reliable baseline. Although SVO at a social angle of 90° achieves slightly lower returns than IPPO in most environments, its angle parameter provides a useful mechanism to control agent behavior.
Table 2: Returns Across Different Environments as a Function of $\theta$ under the SVO Algorithm.
Environment $0^{\circ}$ (Individualistic) $22.5^{\circ}$ $45^{\circ}$ $67.5^{\circ}$ $90^{\circ}$ (Altruistic)
Coins $11.81$ $60.38$ $160.43$ $162.25$ 162.46
Harvest: Open $79.71$ $68.96$ $68.28$ $238.40$ 254.54
Harvest: Closed $81.05$ $74.74$ $74.68$ $298.97$ 309.94
Harvest: Partnership $77.44$ $72.45$ $71.58$ $225.94$ 234.30
Clean Up $0.02$ $0.06$ $50.58$ $1060.25$ 1410.53
Coop Mining $210.26$ $207.96$ $415.99$ $630.32$ 647.61
Mushrooms $5.94$ $51.26$ $291.55$ $321.22$ 400.85
Gift Refinement $104.29$ $105.09$ $105.22$ $107.38$ 227.95
Prisoners Dilemma: Arena $22.67$ $23.27$ $24.13$ $34.96$ 53.36
(a) Coins
(b) Commons Harvest: Open
(c) Commons Harvest: Closed
(d) Commons Harvest: Partnership
(e) Clean Up
(f) Coop Mining
(g) Mushrooms
(h) Gift Refinement
(i) Prisoner Dilemma: Arena
@@ 4.3 Social Dilemma Environments Attributes
Coins (Figure 3(a)), Harvest: Open (Figure 3(b)), Harvest: Closed (Figure 3(c)), Harvest: Partnership (Figure 3(d)) Gift Refinement (Figure 3(h)), Mushrooms (Figure 3(g)) and exhibit fear: namely that agents would prefer to defect if some of their co-players are defecting. An agent actually does better cooperating if all their co-players are cooperating. In Harvest: Open, we observe that it only takes a single defector agent to have a severely negative impact on the returns of other players. This is because defector agents typically harvest the last apple in a patch, permanently depleting it. The cooperator policies have learned to leave a greater amount of apples in each patch, which causes a higher regrowth rate, so they generally exhibit a greater tendency to abstain from harvesting in the presence of a defector. In Harvest: Closed, we find that more than one defector is necessary to inflict a severe negative impact on the other agents. This occurs because Harvest: Closed partitions resources into two distinct rooms. After a lone defector depletes all apples in one room, re-entering the other requires visually locating its entrance and traversing through it.
In Clean Up (Figure 3(e)), we instead observe greed, as there is an incentive to defect if a sufficient number of co-players are cooperating. This is because cooperators will clean the pollution, whereas the defector will not, but it requires a minimum of three cooperators to clean the river for apples to grow. Coop Mining (Figure 3(f)) and Prisoners Dilemma: Arena (Figure 3(i)) also primarily exhibit greed. In these environments, if a sufficient number of agents cooperate, other agents can obtain a higher payoff by defecting. In Coop Mining, this is because cooperators tend to neglect the iron ore, allowing the defector to collect more iron ore. In Prisoner’s Dilemma: Arena, when most agents collect cooperative resources, defectors can gain higher returns through interactions with cooperators.
Metrics Analysis The proposed metrics quantify changes in agent behavior or environment states triggered by cooperative actions. A larger metric (Table 3) value reflects a higher level of cooperative behavior exhibited by the agents. In general, a higher degree of cooperative behavior among agents leads to larger collective returns. However, there are exceptions. In the three Harvest environments, the SVO algorithm encourages altruistic actions, which can reduce the agents’ efficiency in collecting apples. Specifically, agents may intentionally avoid picking apples to allow others access to them. As a result, more apples remain on the map, leading to a decrease in harvest efficiency and ultimately a reduction in collective returns. In more extreme cases, agents may not collect any apples at all, causing the Harvest metric to remain at a high value. Under typical conditions where agents receive rewards based on apple collection, this metric performs well. As expected, IPPO-IR and SVO-$0^{\circ}$, which encourage selfish behavior, produce significantly lower metric values than more cooperative approaches such as IPPO-CR, MAPPO and SVO-$90^{\circ}$.
Table 3: Metrics for Assessing Cooperative and Competitive Tendencies Across Environments.
Envs IPPO-IR IPPO-CR MAPPO SVO-$0^{\circ}$ SVO-$90^{\circ}$
Coins $100.83$ $164.07$ $126.23$ $20.44$ $\mathbf{164.56}$
Harvest: Open $0.91$ $23.68$ $18.45$ $2.97$ $\mathbf{37.67}$
Harvest: Closed $1.71$ $17.57$ $18.57$ $4.61$ $\mathbf{22.70}$
Harvest: Partnership $1.42$ $17.75$ $15.56$ $2.25$ $\mathbf{25.12}$
Clean Up $9.28$ $109.54$ $\mathbf{133.99}$ $18.55$ $99.08$
Coop Mining $298.21$ $521.56$ $\mathbf{571.25}$ $14.26$ $483.24$
Mushrooms $4.07$ $19.55$ $\mathbf{22.96}$ $0.02$ $12.33$
Gift Refinement $0.03$ $\mathbf{28.37}$ $20.24$ $0.03$ $3.83$
Prisoners Dilemma: Arena $58.65$ $73.67$ $\mathbf{119.63}$ $0.10$ $76.24$